OpenStudy (loser66):

A mass weighing 3lb stretches a spring 3 in. If the mass is pushed upward, contracting the spring a distance of 1 in, and then set in motion with a downward velocity of 2ft/s, and if there is no damping, find the position u of the mass at any time t

OpenStudy (loser66):

I have to apply formula \[mU'' +\gamma U' + kU=F(t)\]

OpenStudy (amistre64):

its just a second order differential equation

OpenStudy (amistre64):

let u = e^(rx), u' = re^(rx), u'' = r^2e^(rx) e^(rx) (mr^2 + gr +k) = 0 is the homogenous form r = (-g +- sqrt(g^2-4mk))/(2m) which i believe goes into sin/cos stuff

OpenStudy (amistre64):

we have initial conditions to apply to it to solve for the arbitrary constants